OpenStudy (anonymous):

Use the function: f(x) = x3 + 2 Solving for the inverse will yield: Select one:

OpenStudy (anonymous):

Select one: a. No solution as there is no answer

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

And what would that give me

OpenStudy (anonymous):

Do it...

OpenStudy (anonymous):

I meant what will it look like after I do all that

OpenStudy (anonymous):

@adjax can you help me please

OpenStudy (anonymous):

I want you to take a look at this: http://mathinsight.org/inverse_function_examples y

OpenStudy (anonymous):

do I have to read the whole entire thing

OpenStudy (anonymous):

Follow my steps \[f(x)=x^3+2 \implies y = x^3+2\]

OpenStudy (anonymous):

im following

OpenStudy (anonymous):

\[x=y^3+2\] switch x's and y's now try the next step, which is solving for y.

OpenStudy (anonymous):

how do I solve for y?

OpenStudy (anonymous):

Get y on its own, some algebra, what would your first step be?

OpenStudy (anonymous):

would I subtract 2 from both side

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[y^3=x-2\]

OpenStudy (anonymous):

Next step?

OpenStudy (anonymous):

\[\frac{ x }{ 2 }=y ^{3}\]

OpenStudy (anonymous):

is that what s supposed to look like

OpenStudy (anonymous):

No you subtract from both sides, not divide :p

OpenStudy (anonymous):

\[x=y3+2 \] -2 -2

OpenStudy (anonymous):

don't they both equal 0

OpenStudy (anonymous):

On the right side it gets cancelled out.

OpenStudy (anonymous):

This is odd, I've seen you ask much tougher problems involving higher algebra.

OpenStudy (anonymous):